OpenStudy (ally_b):

Help?? Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively.

OpenStudy (welshfella):

2nd term = a1 r 5th term = a1 r^4 where r is the common ratio and a1 = first term so 5th/2nd term = a1 r^4 / a1 r = r^3 = 768/-12 = = -64 so r^3 = -64

OpenStudy (welshfella):

now you can find the common ratio r r = cube root of -64

OpenStudy (welshfella):

can you work that out? use ur calculator

OpenStudy (ally_b):

yes thank you!

OpenStudy (welshfella):

yw first term will be -12/-4 nth term = a1 r^(n-1)

OpenStudy (ally_b):

So the answer would be 3*(-4)^n-1 @welshfella